# Deep learning joint extremes of metocean variables using the SPAR model

### v1

@@ 4.4 Diagnostic plots
As the SPAR model does not model the tails of the marginal variable directly, it is also useful to assess how well the simulated data matches the observed marginal tails. Quantile-quantile (QQ) plots of the observed and simulated marginal variables are shown along the diagonal in Figure 2. The agreement appears good. However, it is difficult to assess the quality of fit on this scale. Instead, Figure 9 shows exceedance and non-exceedance probabilities on a logarithmic scale for the five variables used in the model, as well as the derived variables $H_{s}$, $U_{10},$ and $s$. Overall, the model performs well and provides a good match for both the upper and lower observed tails. There is a slight tendency to underestimate the upper tail of $H_{x}$, which results in a slight underestimation of the upper tail of $H_{s}$. The over-prediction of steep waves mentioned above is also evident.
@@
Finally, to give an indication of the variation in model performance over the angular domain, Figure 10 shows QQ plots of simulated against observed threshold exceedances, binned over small angular ranges. As above, we have used a grid of 1002 fixed angles and a radius of $15^{\circ}$ to define the angular bins. Only bins with 200 or more observations have been used, so that there are approximately 20 or more threshold exceedances per bin. The local threshold is taken as the average of the empirical 90th percentile of observed and simulated values in each bin (this approximates an average of the non-stationary threshold over the bin). There is some scatter between the observed and simulated values. However, the aggregated trend shows good agreement, although with a slight tendency to underestimate at larger values. This may be related to the use of maximum likelihood estimation, which is known to produce a slight negative bias in quantile estimates for small sample sizes [hosking1987parameter].
Figure 10: QQ plots of simulated vs observed threshold exceedances in bins of $15^{\circ}$ radius from a grid of points on $\mathbb{S}^{4}$. Grey lines are individual bins; red line is aggregated value over all bins; black dashed line is equality.
@@ 5 Discussion and conclusions
In this work, we have introduced a deep learning framework for inference with the SPAR model. The computational scalability and robustness of neural networks result in a modelling approach which requires very few assumptions, offers a high degree of flexibility, and can be applied in higher-dimensional settings compared to existing techniques. We use our approach to approximate the complex joint tail behaviour of a five dimensional metocean dataset, with diagnostics indicating our model is able to accurately represent the observed dependence structure. Given the complex dependence structures observed in the data, the MLP model for the angular variation of the GP parameters performs very well in representing the ‘extremes’ of the dataset. By ‘extremes’, we are referring not just to the largest and smallest values of each observed variable, but anywhere on the outer part of the data cloud. Moreover, the SPAR model provides an asymptotically justified basis for extrapolating outside the range of observations. Simulation from the fitted SPAR model subsequently allows ones to generate large, physically-realistic event sets, allowing practitioners to easily perform robust risk assessments and estimate probabilities of structural failure.
@@
We note that selecting an architecture for any neural network is non-trivial, and care must be taken to ensure the resulting model offers sufficient flexibility without overfitting. We also remark that, in general, neural networks require a large amount of data for accurate model fitting [goodfellow2016deep]; this is generally a challenge for modelling extremes since, by definition, very little data are available. It is currently not clear what sample sizes, tuning parameters, or architectures are required to accurately fit the SPAR model via deep learning, and this should be explored in further work. In the present approach, we have used 80% of the data for training and 20% for validation. The validation data are used to avoid overfitting. It is possible that using only 20% of observations for validation is insufficient to force a sufficiently smooth solution in regions of sparse angular observations. An alternative is to use a full cross-validation scheme, in which the model results are averaged over e.g., five fits using a different 20% of the data for testing. This would make better use of the limited observations, although with increased computational cost.
@@
One notable observation from Section 4 was that the choice of origin for defining the SPAR model was non-trivial. Initially, we naively assumed that the componentwise mean was a suitable origin, but this resulted in issues when fitting the model, as the data cloud was not star-shaped with respect to this initial choice of origin. Selecting an appropriate origin for the SPAR model beyond the lower dimensional ($d\leq 3$) setting remains a challenge as full data visualisation is not possible, and inappropriate choices can invalidate the modelling assumptions. Future work could explore the robust selection of the origin in a data-driven manner, removing the need for domain specific knowledge.
@@
From Section 4, it was also clear that as the dimension of the data increases, so too does the sparsity. This is particularly apparent in Figure 6, where we observed some orthants containing just a handful of observations. Inference within such regions is problematic, since there is very little data to train the neural network model. We also noted that in sparse regions, the resulting parameter estimates did not always respect the physical features of the data, e.g., upper bounds of the variables. Such issues are not unique to the SPAR model, and one would expect to encounter the same problems with alternative modelling approaches when applied in high dimensions. Future work could explore whether notions of sparsity can be incorporated into the SPAR framework to improve the robustness and efficiency of the model.
@@ Acknowledgment
EM was funded by the EPSRC Supergen Offshore Renewable Energy Hub, United Kingdom [grant no: EP/Y016297/1].

### v2

@@ 4.5 Assessment of conditional radial model
As the SPAR model does not model the tails of the marginal variable directly, it is also useful to assess how well the simulated data matches the observed marginal tails. Figure 13 shows exceedance and non-exceedance probabilities on a logarithmic scale for the five variables used in the model, as well as the derived variables $H_{s}$, $U_{10},$ and $s$. The model predicted some unrealistically steep conditions, so we have removed data points where steepness exceeded a value of 0.1. This corresponded to approximately 0.01% of the simulated sample and small wave heights ($H_{s}<3$ m). With these points removed, overall, the model performs well and provides a good match for both the upper and lower observed tails. There is a slight tendency to underestimate the upper tail of $H_{x}$, which results in a slight underestimation of the upper tail of $H_{s}$.
Figure 12: Comparison of empirical joint densities of pairs of variables (coloured plots) with contours of the joint densities from a sample from the SPAR model of 100 times the original sample size (red lines). Note that the colour resolution has been restricted so that the boundaries of the coloured regions correspond to the contour levels from the SPAR model. Contours at equal logarithmic increments.
@@
Finally, to give an indication of the variation in model performance over the angular domain, Figure 14 shows QQ plots of simulated against observed threshold exceedances, binned over small angular ranges. As above, we have used a grid of 1002 fixed angles and a radius of $15^{\circ}$ to define the angular bins. Only bins with 200 or more observations have been used, so that there are approximately 20 or more threshold exceedances per bin. To account for the variation of the threshold level over the bins, the threshold is taken as the empirical quantile at exceedance level $\zeta$ from the simulated dataset (this approximates an average of the non-stationary threshold over the bin). The simulated data are then interpolated to the same probability levels as the observations. There is some scatter between the observed and simulated values. However, the aggregated trend shows good agreement, although with a slight tendency to underestimate at larger values. Although the total sample size is relatively large, the number of observations in local angular regions can be quite small. The underestimation from the model may therefore be related to the use of maximum likelihood estimation, which is known to produce a slight negative bias in quantile estimates for small sample sizes [hosking1987parameter].
Figure 14: QQ plots of simulated versus observed threshold exceedances in bins of $15^{\circ}$ radius from a grid of points on $\mathbb{S}^{4}$. Grey lines are individual bins; red line is aggregated value over all bins; black dashed line is equality.
@@ 5 Discussion and conclusions
In this work, we have introduced a deep learning framework for inference with the SPAR model. The computational scalability and robustness of neural networks result in a modelling approach which requires very few assumptions, offers a high degree of flexibility, and can be applied in higher-dimensional settings compared to existing techniques. We use our approach to approximate the complex joint tail behaviour of a five dimensional metocean dataset, with diagnostics indicating our model is able to accurately represent the observed dependence structure. Given the complex dependence observed in the data, the MLP model for the angular variation of the GP parameters performs very well in representing the ‘extremes’ of the dataset. By ‘extremes’, we are referring not just to the largest and smallest values of each observed variable, but anywhere on the outer part of the data cloud. Moreover, the SPAR model provides an asymptotically justified basis for extrapolating outside the range of observations. Simulation from the fitted SPAR model subsequently allows ones to generate large, physically-realistic event sets, allowing practitioners to easily perform robust risk assessments and estimate probabilities of structural failure.
@@
We note that selecting an architecture for any neural network is non-trivial, and care must be taken to ensure the resulting model offers sufficient flexibility without overfitting. We also remark that, in general, neural networks require a large amount of data for accurate model fitting [goodfellow2016deep]; this is generally a challenge for modelling extremes since, by definition, very little data are available. It is currently not clear what sample sizes, tuning parameters, or architectures are required to accurately fit the SPAR model via deep learning, and this should be explored in further work. In the present approach, we have used 80% of the data for training and 20% for validation. The validation data are used to avoid overfitting. It is possible that using only 20% of observations for validation is insufficient to force a sufficiently smooth solution in regions of sparse angular observations. An alternative is to use a full cross-validation scheme, in which the model results are averaged over, e.g., five fits using a different 20% of the data for testing. This would make better use of the limited observations, although with increased computational cost. Averaging over multiple plausible candidate models, provides a general means of accommodating uncertainty due to modelling choices that are somewhat arbitrary, like the location of the polar origin, the threshold level, and neural network architecture.
@@
One notable observation from Section 4 was that the choice of origin for defining the SPAR model was non-trivial. Initially, we naively assumed that the componentwise mean was a suitable origin, but this resulted in issues when fitting the model, as the data cloud was not star-shaped with respect to this initial choice of origin. Selecting an appropriate origin for the SPAR model beyond the lower dimensional ($d\leq 3$) setting remains a challenge as full data visualisation is not possible, and inappropriate choices can invalidate the modelling assumptions. Future work could explore the robust selection of the origin in a data-driven manner, removing the need for domain-specific knowledge.
In the current work, we have included the wind and wave directions in the model by considering the x- and y-components of the wind speed and wave height. Another way to model the five considered variables is to treat the wind and wave directions as covariates, and then model the joint distribution of wind speed, wave height, and wave period conditional on these two covariates. Initial investigations suggest that this could resolve some of the issues in the current approach caused by the complex shape of the data cloud. If the origin of the polar coordinate system is allowed to vary with wind and wave direction, then the remaining three-dimensional data clouds appear more star-shaped, potentially resulting in a better fit for the model. This will be investigated in future work. Note that if we were to model a two-dimensional case of say wave height and direction, using the SPAR approach, then this is identical to fitting a non-stationary model to $H_{s}$ with wave direction as a covariate. Indeed, this was partially the motivation for the SPAR approach. However, when more variables are involved, treating directions as covariates or considering the joint distribution of x- and y-components does represent different sets of modelling assumptions.
@@
From Section 4, it was also clear that as the dimension of the data increases, so too does the sparsity. This is particularly apparent in Figure 6, where we observed some orthants containing just a handful of observations. Inference within such regions is problematic, since there is very little data to train the neural network model. We also noted that in sparse regions, the resulting parameter estimates did not always respect the physical features of the data, e.g., upper bounds of the variables. Such issues are not unique to the SPAR model, and one would expect to encounter the same problems with alternative modelling approaches when applied in high dimensions. The effect of dimensionality on data sparsity is not necessarily intuitive to understand. For example, generating pseudo-regularly spaced points on $\mathbb{S}^{4}$ in the manner described in Appendix A, with an average spacing of approximately $4.5^{\circ}$ (corresponding to $m=20$ points along the positive half of each dimension), results in 216,002 direction vectors. This corresponds to $216,002/(24\times 365.25)\approx 24.6$ years of hourly observations – a similar order of magnitude to the size of dataset considered here (although, as discussed above, serial correlation in the data reduces the effective sample size). Compare this to the spacing of the same number of direction vectors in two dimensions, which would be $360/216022\approx 0.0017^{\circ}$. So, a dataset that gives a high directional coverage in two dimensions, gives a much more sparse directional coverage in five dimensions. Moreover, the sparsity increases exponentially with the number of dimensions, meaning that this type of analysis (or any multivariate density modelling method) will require an exponential increase in dataset length to maintain a similar level of mean spacing between observed angles. Future work could explore whether notions of sparsity can be incorporated into the SPAR framework to improve the robustness and efficiency of the model.
In the present implementation, the model is only estimated for observations above the threshold level, with the threshold exceedance probability, $\zeta$, chosen by testing models fitted with different values of $\zeta$. A possible extension, would be to model the entire radial distribution with a blended model that takes a semi-parametric form in the bulk of the distribution and a parametric GP form in the tail, similar to [majumder2025semi]. This has the advantage that the uncertainty in threshold estimation is incorporated in the inference.
Finally, we note that although we have accounted for serial correlation in the data in some of the model fitting and diagnostics, the resulting SPAR model is a model for all observations. As such, when we simulate from the model, we do not account for the serial correlation. This could potentially be addressed by using a two-step simulation procedure, where we first simulate a point from the model, then simulate a section of time series which passes through the simulated point. The time series simulation could be accomplished using a resampling approach [mackay2020estimation] or by modelling the temporal evolution of variables conditional on a peak value in a similar manner to [tendijck2023temporal]. This will be investigated in future work.
@@ Acknowledgment
EM was funded by the EPSRC Supergen Offshore Renewable Energy Hub, United Kingdom [grant no: EP/Y016297/1]. This work has made use of the resources provided by the Edinburgh Compute and Data Facility (ECDF) (http://www.ecdf.ed.ac.uk/)
Appendix A Pseudo-regularly spaced points on the hypersphere
To generate a grid of pseudo-regularly spaced points on the hypersphere, we take the approach proposed in [mackay2023-DIFORM], and define a regular grid of points on the $L^{1}$ sphere and project this onto the surface of the $L^{2}$ sphere. This is computed by creating a regular grid of points in the cube $[-1,1]^{d}$, with $12m+$ 1 points along each dimension: $U_{grid}=\{(i_{1}/m,...,i_{d}/m):i_{j}\in\{-m,...,m\},j=1,...,d\}$. Then, we define $U_{1}=\{\mathbf{u}\in U_{grid}:\|\mathbf{u}\|_{1}=1\}$, where $\|\cdot\|_{1}$ is the $L^{1}$ norm, given by $\|\mathbf{u}\|_{1}=|u_{1}|+\cdots+|u_{d}|$ for $\mathbf{u}=(u_{1},...,u_{d})$. Finally, we compute a set of direction vectors $U\subset\mathbb{S}^{d-1}$ by $U=\{\mathbf{u}/\|\mathbf{u}\|_{2}:\mathbf{u}\in U_{1}\}$. This is illustrated in Figure 15 for the case $m=5$ and $d=3$.
Appendix B Simulation from the power spherical distribution
A method for simulating from the PS distribution was presented in [[]Algorithm 1]deCao2020psd. Here we present some background details on simulating from rotationally-symmetric distributions on $\mathbb{S}^{d-1}$, to illustrate why the PS distribution offers computational advantages over the more commonly-used von Mises-Fisher (vMF) distribution. Further details can be found in standard texts on directional statistics, e.g., [mardia2000, ley2017modern]. An essential tool for modelling random vectors on the sphere is the tangent-normal decomposition, illustrated in Figure 16. For any vectors $\mathbf{w},\bm{\mu}\in\mathbb{S}^{d-1}$ we can express $\mathbf{w}$ in terms of components that are aligned to $\bm{\mu}$ and tangential to it:
$\mathbf{w}$ $=t\bm{\mu}+\sqrt{1-t^{2}}\mathbf{v},$
$=\cos(\theta)\bm{\mu}+\sin(\theta)\mathbf{v},$
where $t=\cos(\theta)=\mathbf{w}^{\top}\bm{\mu}$ and $\mathbf{v}=(\mathbf{w}-t\bm{\mu})/\|\mathbf{w}-t\bm{\mu}\|_{2}$. The unit vector $\mathbf{v}$ lies in the tangent space to $\bm{\mu}$, $\mathbb{S}^{d-1}(\bm{\mu}^{\perp})\coloneq\{\bm{x}\in\mathbb{S}^{d-1}:\bm{x}^{\top}\bm{\mu}=0\}$.
Figure 16: Illustration of tangent-normal decomposition of a vector $\mathbf{w}\in\mathbb{S}^{2}$ relative to $\bm{\mu}\in\mathbb{S}^{2}$.
A density $f:\mathbb{S}^{d-1}\to[0,\infty)$ that is rotationally symmetric about its location $\bm{\mu}\in\mathbb{S}^{d-1}$ can be expressed in terms of an angular function $f_{a}:[-1,1]\to[0,\infty)$, as $f(\mathbf{w})=f_{a}(\mathbf{w}^{\top}\bm{\mu})$ for all $\mathbf{w}\in\mathbb{S}^{d-1}$. That is, the density at $\mathbf{w}$ just depends on the distance from $\mathbf{w}$ to $\bm{\mu}$ along the surface of the sphere. The tangent-normal decomposition of a rotationally symmetric random vector leads to the following useful result [watson1983]. Suppose that random vector $\mathbf{W}\in\mathbb{S}^{d-1}$ has density $f$ that is rotationally symmetric about $\bm{\mu}$, with corresponding angular function $f_{a}$. Define the tangent and normal components of $\mathbf{W}$ as above, with $T=\mathbf{W}^{\top}\bm{\mu}$ and $\mathbf{V}=(\mathbf{W}-T\bm{\mu})/\|\mathbf{W}-T\bm{\mu}\|_{2}$. Then $T\in[-1,1]$ and $\mathbf{V}\in\mathbb{S}^{d-1}(\bm{\mu}^{\perp})$ are independent and $\mathbf{V}$ is uniformly distributed in $\mathbb{S}^{d-1}(\bm{\mu}^{\perp})$. Moreover, $T$ has density
$f_{T}(t)=\frac{2\pi^{(d-1)/2}}{\Gamma((d-1)/2)}f_{a}(t)(1-t^{2})^{(d-3)/2}.$ (16)
The distribution of $T$ is referred to as the marginal distribution of $\mathbf{W}$. In words, this result states that we can represent any rotationally symmetric distribution, in terms of one component that describes the distance from the mean direction, and an orientation relative to the mean direction. Due to the rotational symmetry, the distance and orientation are independent.
The tangent-normal decomposition gives a computationally efficient method for simulating from rotationally symmetric distributions [ulrich1984]. First note that if $\bm{e}_{d}=(0,...,0,1)^{\top}$, then $\mathbb{S}^{d-1}(\bm{e}_{d}^{\perp})$ is the embedding of $\mathbb{S}^{d-2}$ into $\mathbb{R}^{d}$ (see Figure 16 - in this example $\mathbb{S}^{2}(\bm{e}_{3}^{\perp})$ is the embedding of the unit circle into $\mathbb{R}^{3}$). Therefore, if we generate a random $T$ from the marginal distribution of $f$ and a uniform random $\mathbf{V}\in\mathbb{S}^{d-2}$, then $\bm{Y}=(T;\sqrt{1-T^{2}}\mathbf{V})$ has density $f$ with location vector $\bm{e}_{1}=(1,0,...,0)^{\top}$. In then remains to rotate $\bm{Y}$ to the appropriate orientation $\bm{\mu}$. This can be achieved using a Householder transformation [householder1958]. The Householder transformation describes a reflection about a (hyper)plane containing the origin with with unit normal vector $\bm{u}$. The transformation can be expressed in terms of the Householder matrix, defined as $\mathbf{P}=\mathbf{I}_{d}-\bm{u}\bm{u}^{T}$, where $\mathbf{I}_{d}$ is the identity matrix and $\bm{u}\bm{u}^{\top}$ is the outer product. For our application, the plane we wish to reflect about has normal vector $\bm{u}=(\bm{e}_{1}-\bm{\mu})/\|\bm{e}_{1}-\bm{\mu}\|_{2}$. Finally, a random vector with mean direction $\bm{\mu}$ is obtained as $\mathbf{W}=\mathbf{P}\bm{Y}$.
Simulation of uniformly distributed points on $\mathbb{S}^{d-2}$ is straightforward. If $N_{1},...,N_{d-1}$ are independent standard normal variables, then $\mathbf{V}=(N_{1},...,N_{d-1})/\|(N_{1},...,N_{d-1})\|_{2}$ is uniformly distributed on $\mathbb{S}^{d-2}$ [muller1959note]. The key feature of the PS distribution that makes it efficient to simulate from, is that the marginal variable $T$ is defined in terms of an affine transformation of a Beta-distributed variable [deCao2020psd]. Namely, $T=2Z-1$, where $Z\sim\text{Beta}(\alpha,\beta)$ with $\alpha=\kappa+(d-1)/2$ and $\beta=(d-1)/2$, where $\kappa$ is the bandwidth parameter (this follows from substituting the expression for the density of the PS distribution (11) into (16)). In contrast, the marginal distribution of the vMF distribution is not related to common distribution types, meaning that rejection sampling is necessary to sample from the marginal, which is computationally inefficient.